OpenStudy (anonymous):

can you help -42=-18-6c?

OpenStudy (shamil98):

isolate c.

OpenStudy (anonymous):

c=2.3

OpenStudy (shamil98):

-42 = -18-6c 42 = 18 + 6c 24 = 6c 4 = c

OpenStudy (anonymous):